INAF queue observing, 2019 January 26-27

Observer: F. Cusano, A. Giunta
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument:

Summary

At the sunset the sky is clear, the wind is at the limit, gusts up to 19 m/s.
A the beginning of the night conditions are not optimal for AO

Night duration (12 degree el twilight) 2:00- 13:10 UT = 11 h 10m

Observing time: 9h 40min (87%)

Weather time loss: 0h *m (%)

technical time loss 1h 30m *(13%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibration Dark LUCI       darks for Nardini q6,q8
Telluric HD18326 LUCI G200 zj+HKspec   luci1.20190127.032-39, luci2.2019017.0007-0016 telluric for Iras 02
Antoniucci Iras 02 LUCI G200 zj+HKspec   luci1.20190127.044-51, luci2.2019017.0021-28  
Telluric HIP44027 LUCI G200+HKspec   luci1.20190127.055-58, luci2.2019017.0032-35 telluric for Marziani
Marziani J085406 LUCI G200+HKspec 0.75h/0.7h luci1.20190127.063-68, luci2.2019017.0040-46 completed
Marziani J093402 LUCI G200+HKspec 0.6h/0.6h luci1.20190127.073-079, luci2.2019017.051-057 completed
Telluric HIP44027 LUCI G200+HKspec   luci1.20190127.083-86, luci2.2019017.0061-64 telluric for Marziani
Telluric HIP56147 LUCI G200+HKspec   luci1.20190127.090-93, luci2.2019017.068-71 telluric for Saracco
Saracco C1-23152 LUCI G200+HKspec 4h/12h luci1.20190127.098-121, luci2.2019017.076-99  
Gilli Target M LUCI G200+HKspec 4h/4h luci1.20190127.126-161, luci2.20190127.104-139 completed
Pinna PX Vir LUCI2 AO     completed ??
Calibration flat, arc LUCI1-2 G200 HKspec     calibration for Gilli

Details

00:00 starting dark for Nardini program, target q6, q8 and telluric
  • luci1.20190127.0001-10 dark for q8
  • luci1.20190127.0011-15 dark for q6
1:31 AO checking, conditions are not optimal for AO observations, seeing unstable and high

1:41 slewing to Cassara' vipD1
  • at 1:44 dimm seeing 2.2
  • 1:45 we wait 5 minutes to see the trend of the seeing
  • 1:55 alignment of the N3.75 on LUCI2
  • 1:57 seeing very variable between 1.5 and 3, conditions are not good for this target we move to backup targets

2:01 we slew to telluric of Antoniucci Iras 02, HD18326
  • 2:02 preset aborted, primary right mirror panic
  • 2:11 dimm seeing 1.5, very variable between 1.2 and 2
  • 2:15 problem solved slewing to telluric
  • acq luci1.20190127.029-31, luci2.2019017.0006-0008
  • sci luci1.20190127.032-39, luci2.2019017.0007-0016
  • 2:29 dimm seeing 2.5, variable between 1.5 and 3

2:31 slewing to Target of Antoniucci Iras 02
  • acq luci1.20190127.040-43, luci2.2019017.0017-20
  • sci luci1.20190127.044-51, luci2.2019017.0021-28
  • 2:43 dimm 1.5

2:53 slewing to telluric of Marziani HIP44027
  • 2:55 two different guide stars, the to check the pointing
  • 3:04 problem in changing the filter manualy, messare Instrument manager busy
  • 3:05 sending again the script, problem with the left side, is not sending any preset
  • 3:07 mos error on LUCI2
  • 3:25 problem solved, sending again the obs
  • acq luci1.20190127.052-54, luci2.2019017.0029-31
  • 3:33 dimm seeing 1.4
  • sci luci1.20190127.055-58, luci2.2019017.0032-35

3:39 slewing to Marziani J085406
  • acq luci1.20190127.059-62, luci2.2019017.0036-39
  • 3:42 dimm seeing 1, seeing is improving
  • sci luci1.20190127.063-68, luci2.2019017.0040-46, from 3:45 to 3:57 problem with camera in LUCI1
  • 3:58 dimm seeing 2.3,again variable now
  • 4:14 dimm seeing 1.3

4:31 slewing to Marziani J093402
  • note same telluric of J085406 so we skip for the moment, we plan to do later another exposure
  • acq luci1.20190127.069-72, luci2.2019017.0047-50
  • 4:34 fwhm on guider 1.5"
  • 4:42 dimm seeing 1.1
  • sci luci1.20190127.073-079, luci2.2019017.051-057
  • 5:08 dimm seeing 0.9

5:10 slewing again to telluric of Marziani HIP44027
  • acq luci1.20190127.080-82, luci2.2019017.058-60
  • 5:24 dimm seeing 0.8
  • sci luci1.20190127.083-86, luci2.2019017.0061-64

5:28 slew to telluric of Saracco HIP56147
  • acq luci1.20190127.087-89, luci2.2019017.065-67
  • 5:32 dimm seeing 0.9
  • sci luci1.20190127.090-93, luci2.2019017.068-71

5:41 slewing to science of Saracco
  • acq luci1.20190127.094-097, luci2.2019017.072-75
  • 5:48 dimm seeing 0.8
  • sci luci1.20190127.098-121, luci2.2019017.076-99
  • 6:26 dimm seeing jumping to 1.4
  • 6:32 dimm seeing back to 0.7
  • 6:57 dimm seeing jump to 1.3 for a couple of minutes, than back to 0.8

8:17 slewing to target M of Gilli
  • check pointing 8:20
  • acq luci1.20190127.122-125, luci2.20190127.100-103
  • 8:33 dimm seeing 1
  • sci luci1.20190127.126-161, luci2.20190127.104-139
  • 9:02 dimm seeing 0.7
  • 10:14 some thin cirrus coming in but the guide star is still ok
  • 10:18 guide star 0.3 mag fainter compared to the beginning now
  • 10:19 dimm seeing 0.7
  • 10:23 now guide star only 0.1 mag fainter
  • 10:33 guide star back to normal

11:04 slew to telluric of Gilli
  • acq luci1.20190127.162-164, luci2.20190127.140-142
  • 11:13 dimm seeing 0.8
  • sci luci1.20190127.165-168, luci2.20190127.143-146

11:20 slewing to Pinna LUCI AO target PX Vir
  • sub-window 128x128 with DIT 0.17
  • 11:25 no guide star found checking pointing
  • 11:26 sending again the script
  • 11:26 dimm seeing 0.7
  • 11:29 alligning the N30 field, huge misalignment, Olga is taking care of it
  • 11:38 alignment done, now clouds are becoming thicker
  • 11:40 guide star not found, clouds, changing the guide star with one brighter
  • 11:44 guide star found
  • 11:50 starting integration
  • 12:00 problem in setting the correct filter
  • 12:32 problem solved starting the obs
  • 12:45 guide star is back to normal flux
  • 13:28 dimm seeing 0.6
  • in 215 and 2016 imgages appear another source or a simple ghost?
  • 13:42 H2 take almost in sun rise

14:37 starting Gilli calibration
  • luci2.20190127.0232-247, luci2.20190127.179-194== flat and arc
Topic revision: r7 - 01 Feb 2019, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback